OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 7 3/4 inches and a width of 4 1/3 inches?

OpenStudy (anonymous):

@myininaya @iambatman @bibby

OpenStudy (anonymous):

@nikato

OpenStudy (calculusfunctions):

Perimeter of a rectangle is the sum of the products of twice the length and twice the width. P = 2L + 2W OR P = 2(L + W)

OpenStudy (anonymous):

?

OpenStudy (nikato):

perimeter= length + length +width +width or 2xlength + 2xwidth

OpenStudy (anonymous):

?

OpenStudy (nikato):

thats the formula for finding the perimeter perimeter= length + length +width +width and you're given length = 7 3/4 and a width = 4 1/3 so plug it in

OpenStudy (anonymous):

ok????????????????????

OpenStudy (nikato):

so whats 7 3/4+ 7 3/4+ 4 1/3+ 4 1/3=?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

1/3

OpenStudy (nikato):

close. its 24. but the fraction part is wrong